OpenStudy (anonymous):

solve equation on the interval [0,2pi) cos^2(x)=cosx dont know what todo, need help where to start please.

myininaya (myininaya):

\[\cos^2(x)-\cos(x)=\cos(x)(\cos(x)-1)=0 => \cos(x)=0 or \cos(x)-1=0\]

OpenStudy (anonymous):

hello myininaya and imran

myininaya (myininaya):

hey

myininaya (myininaya):

i'm so busy i'm not use to this

OpenStudy (anonymous):

site is freezing up

OpenStudy (anonymous):

you can treat this as a regular quadratic equation

OpenStudy (anonymous):

well the answerkey given is in radians,

OpenStudy (anonymous):

but im lost atm,

OpenStudy (anonymous):

yup that right, what you would do is simply: cos^2x-cosx=0 Then factor out a cos and get: cox(x)(cos(x)-1)=0 now set each term to be zero

myininaya (myininaya):

cos(x)=0 when x=pi/2 or 3pi/2 cos(x)-1=0 when cos(x)=1 cos(x)=1 when x=0

OpenStudy (anonymous):

so cosx=0 and cosx=1???

OpenStudy (anonymous):

yup that right

OpenStudy (anonymous):

answer is {pi/4, 3pi/4, 5pi/4, & 7pi/4} i think i use the unit circle right?

OpenStudy (anonymous):

since its restriction is (o,2pi)

OpenStudy (zarkon):

myininaya has the correct answer

myininaya (myininaya):

how did you get that solution set?

OpenStudy (anonymous):

yes, so within that interval you determine when cos x is 0 and 1

OpenStudy (anonymous):

within the interval [0,pi], myin has it right

myininaya (myininaya):

within the interval [0,2pi) you mean

OpenStudy (anonymous):

oh ok thank you, im starting to see it

OpenStudy (anonymous):

yes, sorry i meant [0,pi). Thank you for the correction

myininaya (myininaya):

lol 2pi*

OpenStudy (anonymous):

lol

myininaya (myininaya):

its okay lagrange we know what you mean

OpenStudy (anonymous):

can you tell i am having a horrible day, lol

OpenStudy (anonymous):

OOPPS wrong asnwer key....lol its (0,pi/2, and 3pi/2) sorry now i get it haha